OpenStudy (anonymous):

use half angle identities to solve sin(pi/12), help?

OpenStudy (i0iz0gangster):

it becomes pi/6 because the pi/12 is the half angle and we need to substitute the regular angle which is 2x the half angle. sin pi/12 = + or - sqrt((1-sqrt(3)/2)/2) sin pi/12 = + or - sqrt((2-sqrt(3)/4) That is the answer

OpenStudy (anonymous):

do you have to multiply always by 2?

OpenStudy (anonymous):

sin a/2 = + or - sqrt((1-cosa)/2) sin pi/12 = + or - sqrt((1-cos pi/6)/2) ........ it becomes pi/6 because the pi/12 is the half angle and we need to substitute the regular angle which is 2x the half angle. sin pi/12 = + or - sqrt((1-sqrt(3)/2)/2) sin pi/12 = + or - sqrt((2-sqrt(3)/4) That is the answer (u can find tht on google just saying lool x;